# Computation of the expectation value of the spin operator $\hat{S}^{2}$ for the Spin-Flip Bethe-Salpeter Equation

### v1

@@ 1 Introduction
Spin-flip methods have been introduced in the electronic structure literature beginning in 2001 [1], developed for post-Hartree-Fock methods [2] and time-dependent density-functional theory (TDDFT) [3]. These approaches describe states of an open shell system via spin-flipping excitations from a high-spin reference state that can be accurately described by Hartree-Fock or density-functional theory (DFT). An important aspect of these methods is the evaluation of matrix elements of the spin operator $\hat{S}^{2}$ to assist with analysis of the computed results, and assess the presence of spin contamination [4] which is notorious in such approaches [5]. Ref. [6] presents in its appendix the necessary equation to calculate $\hat{S}^{2}$ for spin-flip excitations. However, a detailed derivation is missing and the authors are not aware of its presence elsewhere in the literature. Obtaining the end result, beginning from the so-called “super-operator” approach, is not trivial. In this article, we provide a thorough derivation of this equation, to complement the derivation of the spin-conserving version presented in Ref. [7]. We include many steps in working out the normal ordering of the creation and annihilation operators in the nested commutators to assist any readers interested in working through the results presented in Ref. [6] and [7], or applying the “super-operator” approach to previously unconsidered transitions. We additionally include information about the practical implementation for calculating matrix elements of $\hat{S}^{2}$ with spin-flip excited states, using examples first considered in Ref. [8] to illustrate results, considerations, and time performance of the relatively straightforward serial Python scripts.
@@
This article is organized as follows. In Section 2, we discuss details regarding our previous SF-BSE calculation results for the test systems, ethylene under torsion and the ${}^{-}$ center in diamond. In Section 3, we discuss why the explicit calculation of $\langle\hat{S}^{2}\rangle$ is necessary. In Section 4, we discuss sources of deviation of the computed value of $\langle\hat{S}^{2}\rangle$ from naive expectation. In Section 5, we provide the derivation of the formula for $\langle\hat{S}^{2}\rangle$ for spin-flip excitations. In Section 6, we use a simple example to illustrate how to use that formula. Finally, in Section 7, we discuss the scripts and their performance in calculating $\langle\hat{S}^{2}\rangle$.
@@ 2 Summary of example calculations with SF-BSE
[…]
@@
The DFT calculation in Ref. [8] of the ethylene orbitals was performed with the DFT code Octopus [17, 18]. We used the Optimized Norm-Conserving Vanderbilt pseudopotentials [19] from the Pseudo-Dojo pseudopotential database [20] with the PBE exchange-correlation [21]. The relaxed atomic coordinates (at $0^{\circ}$) were calculated with a 0.115 Å real-space grid spacing, roughly equivalent to a 115 Ry planewave wavefunction cutoff, in a box with edge-length 12 Å. A smaller box size was used in the subsequent calculations, which contains 99% of the charge density for both the ethylene molecule with no torsion and $90^{\circ}$ of torsion. The ground- and excited-state energies were calculated explicitly at torsions of 0, 5, 10, 15, 30, 45, 60, 75, 80, 85, and ${}^{\circ}$. For DFT input to SF-BSE calculations, we used a less stringent real-space grid spacing for both the wavefunctions and density of 0.34 Bohr, equivalent to about 85 Ry planewave cutoff for the wavefunctions. We obtained the $M_{S}=1$ reference state by constraining occupations. Based on [22], we used 860 empty states and 24 Ry for the calculation of the screened Coulomb interaction. For the SF-BSE Hamiltonian, 5 occupied spin-up (the maximum allowed) and 55 unoccupied spin-down orbitals were used.
@@
The ${}^{-}$ center in diamond is a well-studied open-shell defect in a crystalline solid. Within its bulk band gap are so-called in-gap single-particle states labeled “$v$” and “${e_{x},e_{y}}$”, the latter being doubly-degenerate. (See, for instance, Ref. [23].) Its ground state is a triplet, with the $S=1,M_{S}=1$ state having its up-spin $v$, $e_{x}$, and $e_{y}$ orbitals occupied and only its down-spin $\bar{v}$ orbital occupied. (For SF-BSE calculations, this is our high-spin reference state.) Its three-fold degenerate triplet ground state is labeled ${}^{3}A_{2}$, and its optical excitations, therefore, are not accessible by conventional GW/BSE. We compute its transition energies to the singlet states ${}^{1}E$, ${}^{1}A_{1}$ and the triplet state ${}^{3}E$ in Ref. [8]. Again, we compute the bands in DFT with the code Octopus [17, 18]. The defect is placed in a 2$\times$2$\times$2 supercell, with 63 atoms (62 carbon, 1 nitrogen), since the orbitals of the in-gap states are well-localized [23]. The real-space grid spacing is 0.34 Bohr, again, equivalent to an 85 Ry planewave cutoff. The pseudopotentials are also Optimized Norm-Conserving Vanderbilt pseudopotentials [19] from the Pseudo-Dojo pseudopotential database [20] with the PBE exchange-correlation[21]. The dielectric matrix is calculated with 300 empty states and a cutoff of 12 Ry, consistent with the choice of parameters in Ref. [24]. The SF-BSE Hamiltonian is constructed from 12 occupied and 11 unoccupied orbitals.
@@ 3 Necessity for the explicit calculation of $\langle\hat{S}^{2}\rangle$
Let us consider an example: the excitations as computed from SF-BSE for ethylene, at zero torsion. The single-particle orbitals that describe the HOMO and LUMO (as a minimal basis set) are $|\pi\uparrow\rangle$, $*|\pi^{*}\uparrow\rangle$, $|\bar{\pi}\downarrow\rangle$, and $*|\bar{\pi}^{*}\downarrow\rangle$. The use of the overbars for the down-spin states denotes the different spatial wavefunction, compared to the up-spin state, as occurs in a spin-polarized or spin-unrestricted approach. The high-spin reference state is the triplet $*|\pi\uparrow,\pi^{*}\uparrow\rangle$. The excitation vectors of interest describe the following states[25], with approximately the following numerical coefficients (to two decimals) and signs[8]:
@@
$|Z\rangle$ $*\displaystyle=-0.20|\pi\uparrow,\bar{\pi}\downarrow\rangle+0.98|\pi^{*}\uparrow,\bar{\pi}^{*}\downarrow\rangle\,$ (6)
In a “spin-restricted” calculation, the eigenstates from a spin-unpolarized calculation (i.e., disregarding spin entirely), are doubled and placed into two spin channels, with the occupations made consistent with the magnetization of the system, post facto. The orbitals in both spin channels are necessarily the same, including the energies and phase information. However, in spin-polarized calculations, in which the different spin channels have independent Hamiltonians, the orbitals with the same band or orbital index but different spins generally have different real-space wavefunctions and energies. (These are called “spin-unrestricted calculations”). In a spin-unrestricted calculation, therefore, there will likely be some phase rotation of the orbitals in the down-spin channel relative to the up-spin channel. If, for example, $*|\bar{\pi}^{*}\downarrow\rangle$ orbital has a change relative to the $*|\pi^{*}\uparrow\rangle$ orbital, a naive reading of the excitation vector coefficients would swap the interpretation of the $|T\rangle$ and $|S\rangle$ states.
@@
Additionally, the particular states one is interested in describing may be most quickly identified by their $\langle\hat{S}^{2}\rangle$ values. For instance, the $Z$ state of ethylene, under zero torsion, is the tenth excited state, with several maximally contaminated (i.e., equal parts singlet and triplet) uninteresting states intermediate in energy between it and the lower-energy excited singlet $V$ state. While this state at zero torsion may be identified by reading the coefficients for several excitation eigenvectors, at torsion angles greater for zero, this laborious brute-force approach will not be productive. In fact, the order of the $Z$ state changes as a function of torsion angle, ultimately becoming at $90^{\circ}$ of torsion the fourth excitation as ordered by energy.
Figure 2: The frontier orbitals of ethylene under zero torsion, in the high-spin reference state (up-spin orbitals occupied). The down-spin $\bar{\pi}^{*}$ orbital in this illustration has a reversal of phase relative to the up-spin orbital.
@@ 4 Spin Contamination
The eigenvectors computed from SF-BSE may yield an expectation value of $\hat{S}^{2}$ that differs from proper values of $S\left(S+1\right)$ such as 0 (singlet, $S=0$) or 2 (triplet $S=1$). One source of spin contamination that even appears in the high-spin reference state before any spin-flip excitation is applied is related to the difference in the spatial wavefunctions for the different spin channels from spin-polarized DFT [26]. In the examples of ethylene under torsion, or the ${}^{-}$, this causes deviations of 0.01 or less. A larger source of spin contamination in SF-BSE comes from the absence of particular transitions from the set of target states that are required to form an eigenvector of $\hat{S}^{2}$ [5]. For example, if the complete state is $\frac{1}{\sqrt{2}}\left(|r\uparrow,\bar{p}\downarrow\rangle+|p\uparrow,\bar{r}\downarrow\rangle\right)$, sometimes one of these requisite target transitions may be inaccessible from a single spin flip from the up-spin channel to the down-spin channel. This leads to maximally-contaminated states, where the computed $\langle\hat{S}^{2}\rangle\approx 1$.
@@ 5 Derivation of $\langle\hat{S}^{2}\rangle$ for spin-flip transitions
Using the Löwdin Formula, which is valid for a many-body wavefunction that is a Slater determinant as in Hartree-Fock, implies approximating $\langle\hat{S}^{2}\rangle_{0}$ via a Slater determinant formed of the KS states of the high-spin reference state [16]. (More sophisticated DFT approximations have also been developed [28, 29].) In Fig. 3, we see that $\langle\hat{S}^{2}\rangle_{0}$ deviates by at most 0.1% from 2.0 for ethylene under any torsion angle, as expected for the triplet spin symmetry of the high-spin reference state. Likewise, $\langle\hat{S}^{2}\rangle_{0}$ for the triplet ground state of the ${}^{-}$ center is 2.05. These results are consistent with a DFT study [30] finding that solids generally have more spin contamination than molecules.
@@
For spin-flip excitations from the high-spin reference state, we compute the difference “$\Delta\langle\hat{S}^{2}\rangle$”between $\langle\hat{S}^{2}\rangle$ for the $I$’th spin-flipped excitation and the reference state: $\Delta\langle\hat{S}^{2}\rangle=\langle\hat{S}^{2}\rangle_{I}-\langle\hat{S}^{2}\rangle_{0}$. While the formula for $\Delta\langle\hat{S}^{2}\rangle$ in the context of SF-TDDFT was previously published as Eq. A10 in [6], was evidently implemented in codes for use in calculations such as [31], and was used for SF-BSE in [11], the derivation does not seem to have been shown in detail and so we work through the derivation here.
@@
$\Delta\Gamma_{r\uparrow,\bar{q}\downarrow,\bar{p}\downarrow,s\uparrow}=\sum_{i\uparrow,\bar{a}\downarrow,j\uparrow,\bar{b}\downarrow}\left(A^{I}_{i\uparrow,\bar{a}\downarrow}\right)^{*}A^{I}_{j\uparrow,\bar{b}\downarrow}\times$
@@
[…]
6 Example application: The defect states of the ${}^{-}$ center
The general equation for $\Delta\langle\hat{S}^{2}\rangle$ is clarified by considering a simple and illustrative example. We will consider a by-hand calculation of Eq. 5 for the $S=1,M_{S}=0$ ground state of the ${}^{-}$ center, with a minimal basis set, and with the same orbitals for the different spin channels (as in a “spin-restricted” calculation).
In a more complete description, the ${}^{-}$ center has four electrons available to occupy six in-gap orbitals [33]: $|v\uparrow\rangle$, $|\bar{v}\downarrow\rangle$ , $|e_{x}\uparrow\rangle$, $|\bar{e}_{x}\downarrow\rangle$ , $|e_{y}\uparrow\rangle$, $|\bar{e}_{y}\downarrow\rangle$, where $e_{x}$ and $e_{y}$ are degenerate single-particle states. The high-spin (triplet) reference state $|v\uparrow,\bar{v}\downarrow,e_{x}\uparrow,e_{y}\uparrow\rangle$ has $S=1,M_{S}=1$. From this high-spin reference state, we find from actual SF-BSE calculations the $S=1,M_{S}=0$ triplet state

### v2

@@ 1 Introduction
Spin-flip methods have been introduced in the electronic structure literature beginning in 2001 [1], developed for post-Hartree-Fock methods [2] and time-dependent density-functional theory (TDDFT) [3]. These approaches describe states of an open shell system via spin-flipping excitations from a high-spin reference state that can be accurately described by Hartree-Fock or density-functional theory (DFT). The approach was initially used for open-shell molecules[1], but recently SF-TDDFT has been successfully applied to defects in solids [4]. An important aspect of these methods is the evaluation of matrix elements of the spin operator $\hat{S}^{2}$ to assist with analysis of the computed results, and assess the presence of spin contamination [5] which is notorious in such approaches [6]. While Ref. [4] does not comment on the spin contamination of their computed results for energy levels of defected solids, analysis of the expectation value of $S^{2}$ for these states would be useful. Ref. [7] presents in its appendix the necessary equation to calculate $\hat{S}^{2}$ for spin-flip excitations. However, a detailed derivation is missing and the authors are not aware of its presence elsewhere in the literature. Obtaining the end result, beginning from the so-called “super-operator” approach, is not trivial. In this article, we provide a thorough derivation of this equation, to complement the derivation of the spin-conserving version presented in Ref. [8]. We include many steps in working out the normal ordering of the creation and annihilation operators in the nested commutators to assist any readers interested in working through the results presented in Ref. [7] and [8], or applying the “super-operator” approach to previously unconsidered transitions. We additionally include information about the practical implementation for calculating matrix elements of $\hat{S}^{2}$ with spin-flip excited states, using examples first considered in Ref. [9] to illustrate results, considerations, and time performance of the relatively straightforward serial Python scripts.
@@
This article is organized as follows. In Section 2, we discuss details regarding our previous SF-BSE calculation results for the test systems, ethylene under torsion and the NV- center in diamond. In Section 3, we discuss why the explicit calculation of $\langle\hat{S}^{2}\rangle$ is necessary. In Section 4, we discuss sources of deviation of the computed value of $\langle\hat{S}^{2}\rangle$ from naive expectation. In Section 5, we provide the derivation of the formula for $\langle\hat{S}^{2}\rangle$ for spin-flip excitations. In Section 6, we use a simple example to illustrate how to use that formula. Finally, in Section 7, we discuss the scripts and their performance in calculating $\langle\hat{S}^{2}\rangle$.
@@ 2 Summary of example calculations with SF-BSE
The authors have previously calculated the ground and excited states for interesting open-shell systems in Ref. [9], using an implementation with a modified version of the BerkeleyGW code [15], revision 7294, roughly equivalent to public release version 2.1. Among these, first, is the ethylene molecule (C2H4) under torsion (from $0^{\circ}$ to $90^{\circ}$). The second is the NV- center in diamond, a defect in crystalline diamond-structure carbon in which one carbon atom is substituted with nitrogen (“N”), a neighboring atom is left vacant (“V”), and the defect bears an overall negative charge (“-”). Here we summarize the key states and computational parameters for these previous calculations, which we further analyze in this work.
@@
The DFT calculation in Ref. [9] of the ethylene orbitals was performed with the DFT code Octopus [18, 19], version 8.4. We used the Optimized Norm-Conserving Vanderbilt pseudopotentials [20] from the Pseudo-Dojo pseudopotential database [21], version 0.2, with the PBE exchange-correlation [22]. The relaxed atomic coordinates (at $0^{\circ}$) were calculated with a 0.115 Å real-space grid spacing, roughly equivalent to a 115 Ry planewave wavefunction cutoff, in a box with edge-length 12 Å. A smaller box size was used in the subsequent calculations, which contains 99% of the charge density for both the ethylene molecule with no torsion and $90^{\circ}$ of torsion. The ground- and excited-state energies were calculated explicitly at torsions of 0, 5, 10, 15, 30, 45, 60, 75, 80, 85, and 90. For DFT input to SF-BSE calculations, we used a less stringent real-space grid spacing for both the wavefunctions and density of 0.18 Å , equivalent to about 85 Ry planewave cutoff for the wavefunctions. We obtained the $M_{S}=1$ reference state by constraining occupations. Based on [23], we used 860 empty states and 24 Ry for the calculation of the screened Coulomb interaction. For the SF-BSE Hamiltonian, 5 occupied spin-up (the maximum allowed) and 55 unoccupied spin-down orbitals were used.
@@
The NV- center in diamond is a well-studied open-shell defect in a crystalline solid. Within its bulk band gap are so-called in-gap single-particle states labeled “$v$” and “${e_{x},e_{y}}$”, the latter being doubly-degenerate. (See, for instance, Ref. [24].) Its ground state is a triplet, with the $S=1,M_{S}=1$ state having its up-spin $v$, $e_{x}$, and $e_{y}$ orbitals occupied and only its down-spin $\bar{v}$ orbital occupied. (For SF-BSE calculations, this is our high-spin reference state.) Its three-fold degenerate triplet ground state is labeled ${}^{3}A_{2}$, and its optical excitations, therefore, are not accessible by conventional GW/BSE. We compute its transition energies to the singlet states ${}^{1}E$, ${}^{1}A_{1}$ and the triplet state ${}^{3}E$ in Ref. [9]. Again, we compute the bands in DFT with the code Octopus [18, 19]. The defect is placed in a 2$\times$2$\times$2 supercell, with 63 atoms (62 carbon, 1 nitrogen), since the orbitals of the in-gap states are well-localized [24]. The real-space grid spacing is 0.34 Bohr, again, equivalent to an 85 Ry planewave cutoff. The pseudopotentials are also Optimized Norm-Conserving Vanderbilt pseudopotentials [20] from the Pseudo-Dojo pseudopotential database [21] with the PBE exchange-correlation[22]. The dielectric matrix is calculated with 300 empty states and a cutoff of 12 Ry, consistent with the choice of parameters in Ref. [25]. The SF-BSE Hamiltonian is constructed from 12 occupied and 11 unoccupied orbitals.
@@ 3 Necessity for the explicit calculation of $\langle\hat{S}^{2}\rangle$
Let us consider an example: the excitations as computed from SF-BSE for ethylene, at zero torsion. The single-particle orbitals that describe the HOMO and LUMO (as a minimal basis set) are $|\pi\uparrow\rangle$, $|\pi^{*}\uparrow\rangle$, $|\bar{\pi}\downarrow\rangle$, and $|\bar{\pi}^{*}\downarrow\rangle$. The use of the overbars for the down-spin states denotes the different spatial wavefunction, compared to the up-spin state, as occurs in a spin-polarized or spin-unrestricted approach. The high-spin reference state is the triplet $|\pi\uparrow,\pi^{*}\uparrow\rangle$. The excitation vectors of interest describe the following states[26], with approximately the following numerical coefficients (to two decimals) and signs[9]:
@@
$|Z\rangle$ $=-0.20|\pi\uparrow,\bar{\pi}\downarrow\rangle+0.98|\pi^{*}\uparrow,\bar{\pi}^{*}\downarrow\rangle\,$ (6)
In a “spin-restricted” calculation, the eigenstates from a spin-unpolarized calculation (i.e., disregarding spin entirely), are doubled and placed into two spin channels, with the occupations made consistent with the magnetization of the system, post facto. The orbitals in both spin channels are necessarily the same, including the energies and phase information. However, in spin-polarized calculations, in which the different spin channels have independent Hamiltonians, the orbitals with the same band or orbital index but different spins generally have different real-space wavefunctions and energies. (These are called “spin-unrestricted calculations”). In a spin-unrestricted calculation, therefore, there will likely be some phase rotation of the orbitals in the down-spin channel relative to the up-spin channel. If, for example, $|\bar{\pi}^{*}\downarrow\rangle$ orbital has a change relative to the $|\pi^{*}\uparrow\rangle$ orbital (see Fig. 2), a naive reading of the excitation vector coefficients would swap the interpretation of the $|T\rangle$ and $|S\rangle$ states.
@@
Additionally, the particular states one is interested in describing may be most quickly identified by their $\langle\hat{S}^{2}\rangle$ values. For instance, the $Z$ state of ethylene, under zero torsion, is the tenth excited state, with several maximally contaminated (i.e., equal parts singlet and triplet) uninteresting states intermediate in energy between it and the lower-energy excited singlet $V$ state. While this state at zero torsion may be identified by reading the coefficients for several excitation eigenvectors, at torsion angles greater for zero, this laborious brute-force approach will not be productive. In fact, the energy ordering of the $Z$ state changes as a function of torsion angle, ultimately becoming at $90^{\circ}$ of torsion the fourth excitation as ordered by energy.
Figure 2: The frontier orbitals of ethylene under zero torsion, in the high-spin reference state (up-spin orbitals occupied), plotted with XCrysDen [27]. The down-spin $\bar{\pi}^{*}$ orbital in this illustration has a reversal of phase relative to the up-spin orbital.
@@ 4 Spin Contamination
The eigenvectors computed from SF-BSE may yield an expectation value of $\hat{S}^{2}$ that differs from proper values of $S\left(S+1\right)$ such as 0 (singlet, $S=0$) or 2 (triplet $S=1$). One source of spin contamination that even appears in the high-spin reference state before any spin-flip excitation is applied is related to the difference in the spatial wavefunctions for the different spin channels from spin-polarized DFT [28]. In the examples of ethylene under torsion, or the NV-, this causes deviations of 0.01 or less. A larger source of spin contamination in SF-BSE comes from the absence of particular transitions from the set of target states that are required to form an eigenvector of $\hat{S}^{2}$ [6]. For example, if the complete state is $\frac{1}{\sqrt{2}}\left(|r\uparrow,\bar{p}\downarrow\rangle+|p\uparrow,\bar{r}\downarrow\rangle\right)$, sometimes one of these requisite target transitions may be inaccessible from a single spin flip from the up-spin channel to the down-spin channel. This leads to maximally-contaminated states, where the computed $\langle\hat{S}^{2}\rangle\approx 1$.
@@ 5 Derivation of $\langle\hat{S}^{2}\rangle$ for spin-flip transitions
Readers interested only in the final result should refer to Eq. 5. The detailed work that follows may assist readers in following the work in, for example, Refs. [7, 17, 8].
@@
Using the Löwdin Formula, which is valid for a many-body wavefunction that is a Slater determinant as in Hartree-Fock, implies approximating $\langle\hat{S}^{2}\rangle_{0}$ via a Slater determinant formed of the KS states of the high-spin reference state [17]. (More sophisticated DFT approximations have also been developed [30, 31].) In Fig. 3, we see that $\langle\hat{S}^{2}\rangle_{0}$ deviates by at most 0.1% from 2.0 for ethylene under any torsion angle, as expected for the triplet spin symmetry of the high-spin reference state. Likewise, $\langle\hat{S}^{2}\rangle_{0}$ for the triplet ground state of the NV- center is 2.05. These results are consistent with a DFT study [32] finding that solids generally have more spin contamination than molecules.
@@
For spin-flip excitations from the high-spin reference state, we compute the difference “$\Delta\langle\hat{S}^{2}\rangle$”between $\langle\hat{S}^{2}\rangle$ for the $I$’th spin-flipped excitation and the reference state: $\Delta\langle\hat{S}^{2}\rangle=\langle\hat{S}^{2}\rangle_{I}-\langle\hat{S}^{2}\rangle_{0}$. (It is standard to subscript $\Delta\langle\hat{S}^{2}\rangle$ with the label for the state $I$ for which it is computed; we will be suppressing this subscript throughout.) While the formula for $\Delta\langle\hat{S}^{2}\rangle$ in the context of SF-TDDFT was previously published as Eq. A10 in [7], was evidently implemented in codes for use in calculations such as [33], and was used for SF-BSE in [12], the derivation does not seem to have been shown in detail and so we work through the derivation here.
@@
$\Delta\Gamma_{r\uparrow,\bar{q}\downarrow,\bar{p}\downarrow,s\uparrow}=\sum_{i\uparrow,\bar{a}\downarrow,j\uparrow,\bar{b}\downarrow}\left(A^{I}_{i\uparrow,\bar{a}\downarrow}\right)^{*}A^{I}_{j\uparrow,\bar{b}\downarrow}\times$ (13)
@@
This equation has an important property of gauge-invariance which can be explicitly verified: a phase change of an individual orbital does not affect the final result, provided that the coefficients $A$ have the counteracting change that keeps the quasiparticle wavefunction $|\Psi^{\textit{I}}\rangle$ constant.
6 Example application: The defect states of the NV- center
The general equation for $\Delta\langle\hat{S}^{2}\rangle$ is clarified by considering a simple and illustrative example. We will consider a by-hand calculation of Eq. 5 for the $S=1,M_{S}=0$ ground state of the NV- center, with a minimal basis set, and with the same orbitals for the different spin channels (as in a “spin-restricted” calculation).
In a more complete description, the NV- center has four electrons available to occupy six in-gap orbitals [35]: $|v\uparrow\rangle$, $|\bar{v}\downarrow\rangle$ , $|e_{x}\uparrow\rangle$, $|\bar{e}_{x}\downarrow\rangle$ , $|e_{y}\uparrow\rangle$, $|\bar{e}_{y}\downarrow\rangle$, where $e_{x}$ and $e_{y}$ are degenerate single-particle states. The high-spin (triplet) reference state $|v\uparrow,\bar{v}\downarrow,e_{x}\uparrow,e_{y}\uparrow\rangle$ has $S=1,M_{S}=1$. From this high-spin reference state, we find from actual SF-BSE calculations the $S=1,M_{S}=0$ triplet state